Eight students in a physics class of twenty students took the final exam and scored an average of 87. What is the average score for the remaining twelve students in the class if the average for the entire class was an 84?
Answer by Boreal(15235)

average=total/n
the total for the class is 84*20=1680
8*87=696
1680-696=984
That is the total for the other 12.
Their average is 984/12=82.
12*-2 below average=-24
8*3 above average=+24, checks